What is the answer for "u" is this equation:
5.1u-6.6=2.4-0.9u
At first I remembered like families so 5.1u and 0.9u will be together I dont know if Im correct so it will be:
5.1u + or - 0.9 I am always getting confused can you explain the steps for me.

Answer by stanbon(75887)

5.1u -6.6 = 2.4 -0.9u
----
Add 0.9u to both sides:
6u - 6.6 = 2.4
-----
Add 6.6 to both sides:
6u = 9
----
Divide both sides by 6 to get:
u = 3/2
================
Cheers,
Stan H.